OpenStudy (immings):

Please Check, getting weird answer. object is thrown in the air [...] modelled by h(t)=-5t^2+20t+1 [...] Find Instantaneous Velocity at t= 2 h(t+h) - h(t) / h

OpenStudy (immings):

h(t) = 21 [plugged in t=2] h(2+h) = -5(2+h)^2 +20(2+h) + 1 (25h^2 +100h+100) + 40 +20h + 1 h(2+h)=25h^2+120h+141 h(2+h)-h(2) / h 25h^2+120h+141 - 21 / h h(25h + 120 +141) -21 /h 25h+240 25(0) + 240 =240m/s

OpenStudy (cruffo):

second line: -5(2+h)^2 +20(2+h) + 1 -5(4 + 4h + h^2) + 40h + 20 + 1 -20 - 20h - 5h^2 + 40h + 20 + 1

OpenStudy (immings):

Thanks

OpenStudy (cruffo):

np